# Combining Static and Dynamic Approaches for Mining and Testing Constraints for RESTful API Testing

### v1

@@ 7.1.1. Methodology
We ran our LLM-based static method to detect constraints. We then manually evaluated each mined constraint.
We used two datasets: 1) the AGORA (Alonso et al., 2023) dataset, and 2) a self-collected dataset (will be explained later). For overlapping analysis, we grouped the invariants from AGORA into the groups for one specific variable, two variables, and so on. We then compared the groups of invariants and constraints on the same set of variables.
As evaluation metrics, we report the number of detected constraints, the number of True Positives (TP), False Positives (FP), and the precision $P$, with $\text{P}=\frac{\text{TP}}{\text{TP}+\text{FP}}$. Specifically, for AGORA, we reuse their experimental results (Total, TP, and P) on their dataset.
7.1.2. Results on the AGORA dataset
Table 1 shows the results on the AGORA dataset with 11 API operations on 7 APIs. Our LLM-based static method identified 217 constraints, with 198 true positives, resulting in a precision of 91.2%. In contrast, AGORA detected 724 true positives out of 989 invariants, resulting in a precision of 73.2%. Notably, a variable has only one constraint, while it can have multiple invariants from AGORA.
RBCTest combines the constraint results from both LLM-based static method and the dynamic approach in AGORA. Thus, we also conducted an overlapping analysis between them. Overlapping constraints are those applied to the same variables, and we compare how well the constraints are detected. A constraint is considered as better than one invariant or a group of invariants if it refers to a narrower set of values for the variable(s) while still adhering to the variables’ description. Similar definition is used for a better invariant or group of invariants. They are equivalent if they cover exactly the same set of possible values for the variable(s).
Our overlapping analysis results reveal that 107 constraints were uniquely detected by our LLM-based static method, while 527 were exclusively identified by AGORA. We further investigated the invariants detected only by AGORA and found that 319 of them pertained to variables lacking descriptions in API specification. This suggests that these invariants were only detectable through the API executions. These invariants primarily involved checks such as 1) whether a variable is URL (33%), 2) a substring of another variable (32%), 3) equal to another variable (13%), or 4) related to string length (7%), and (5) 15.6% covering other types.
Conversely, the constraints uniquely detected by our LLM-based static method mainly occurred in scenarios where the API specification provided variable descriptions, but the AGORA dataset’s API responses did not include these variables (often optional fields). AGORA’s dependency on the diversity of API responses at runtime limits its detection capability in such cases. For instance, in the YouTube API, there are 14 distinct rating schemas that appear only in specific request regions. If AGORA’s API calls do not cover all regions, these schemas remain undetected.
7.1.3. Result Analysis
A closer look at the cases where our LLM-based static method was superior reveals two main reasons for its better performance. First, our LLM-based static method handled more specific domains or ranges of values. For example, in the Amadeus Hotel API, the roomQuantity value is specified as ”an integer between 1 and 9.” Our LLM-based static method correctly identified this constraint and generated an appropriate test, whereas AGORA provided a general invariant ”Numeric,” encompassing any integer, float, etc. Similarly, in the Spotify API, AGORA expected the thumbnailHeight to be ”one of (64, 300, 640),” based on the observed data at runtime, but our LLM-based static method correctly identified it as ”image height in pixels,” which implies any positive integer. Second, our LLM-based static method excelled in mining specific constraints. For instance, in the Amadeus Hotel API, the sellingTotal is defined as ”= Total + margins + markup + totalFees - discounts.” While AGORA simply concluded that sellingTotal was numeric, our LLM-based static method was able to be more specific in the constraint mined from the specification. However, there were a few cases where AGORA outperformed our LLM-based static method. Those cases often involve invariants verifying the format of variables containing URL, and our LLM-based static method sometimes treated URLs as mere strings without further validation.
7.1.4. Combined method
Table 2. RBCTest: Combining constraints and invariants
API-Op Static Dyn. Unq. S Unq. D Dist. Overlapping Combination
Cons. Incons. Total TP P%
A.Hotel 39 61 20 34 54 11 8 109 73 67
GitHub 13 194 3 175 178 4 6 194 188 96.9
GitHub’ 7 127 7 121 128 0 0 152 128 84.2
Marvel 23 55 13 35 48 4 6 113 58 51.3
OMDB 4 15 3 13 16 1 0 18 17 94.4
OMDB’ 1 5 0 3 3 1 0 6 4 66.7
Spotify 21 41 6 35 41 7 3 47 47 100
Spotify’ 12 58 1 29 30 6 5 50 41 82
Spotify” 15 45 1 11 12 6 6 53 43 81.1
Yelp 1 4 1 11 12 0 0 31 12 38.7
YouTube 62 111 45 52 97 5 12 151 114 75.5
Total 198 724 107 527 634 45 46 924 725 78.5
  • Unq: Uniquely detected, Dist: Distinct, (In)Cons: (In)Consistent.
From Table 2, we can see that combining both static and dynamic approaches yield a better result than that of individual method. In total, RBCTest has 107 constraints detected only by our static method, 527 detected only by AGORA. For cases where both approaches detected constraints and invariants on the same variables, in total, we identified 91 overlapping constraints and invariants. Among them, 46 cases are inconsistent. Inconsistencies occur where an invariant extracted from execution data defines a set of instances outside of the set defined by the respective constraints mined from the specification. The inconsistencies detected by RBCTest may reflect a coding bug or out-of-date specification.
As seen in Table 2, RBCTest (the combined method) detects a total of 924 constraints, of which 725 are true positives, i.e., 78.5% precision. RBCTest can identify constraints from both views while maintaining sufficient precision. Despite of lower precision than the LLM-based method, RBCTest detects more true positive constraints (725) in comparison to the individual methods, that is, the 198 true positive constraints from the LLM-based method and the 724 individual invariants and 618 variable-based grouped invariants from AGORA.
7.1.5. Results on the RBCTest dataset
For generalization, we repeated the experiment on our collected dataset (Table 3), having 8 APIs with 83 operations and ‘GET’, ‘POST’, and ‘PUT’. We manually reviewed the API specifications and identified a set of 600 correct constraints as the oracle. Due to some services requiring enterprise subscriptions, we could not execute AGORA on this dataset.
As seen in Table 3, RBCTest via LLM-based static method (no AGORA) identified 457 constraints, corresponding to 457 test cases generated by our LLM-based static method. This includes 28 false positives and 171 missed constraints, yielding an overall precision of 93.9% and a recall of 71.5%, with an F1 score of 81.2%. All standard deviations were below 5%, indicating consistently strong performance across different APIs. The inaccuracy mainly arises from the effects of the filter and verifier: in an effort to minimize the potential for erroneous outputs from the LLM, we restricted both the input to and the output from the model to ensure accuracy. This result is consistent with that in Section 7.1.2.
Our analysis reveals that response-property constraints—which apply to a single response property—are typically straightforward, focusing on format or value range. They account for 32 of the 429 detected constraints. However, in GitLab services, they were difficult to detect due to sparse property descriptions. In contrast, request-response constraints were more prevalent, as they involve multiple variables, capturing complex relationships where a request parameter can influence multiple response properties.
Table 3. Constraints Test Generation on RBCTest Dataset
API # Op. Methods Type GT RBCTest
TP FP FN PRF1%
C.Holiday 4 GET RP 24 16 0 8 100 66.7 80
G.Branch 5 GET,POST RR 49 35 3 14 92.1 71.4 80.4
G.Commit 11 GET,POST,PUT RR 73 54 3 19 94.7 74 83.1
G.Groups 14 GET,POST,PUT RR 85 61 3 24 95.3 71.8 81.9
G.Issues 21 GET,POST,PUT RR 141 92 3 49 96.8 65.2 77.9
G.Project 15 GET,POST,PUT RR 144 110 11 34 90.9 76.4 83
G.Repo. 3 GET,POST RR 44 33 3 11 91.7 75 82.5
Stripe 10 GET,POST RR 19 12 1 7 92.3 63.2 75
RP 21 16 1 5 94.1 76.2 84.2
Total 83 600 429 28 171 93.9 71.5 81.2
Stdev 2.9 4.9 2.9
  • Services: Canada Holidays, GitLab {Branch, Commit, Groups, Issues, Project, Repository}, and Stripe. # of Operations (No. Ops), # of ground truth constraints (GT), Precision (P), and Recall (R). RR for Request-Response Constraints, and RP for Response-Property Constraints.
7.2. Constraint Mining Accuracy (RQ2)
@@
Table 4. LLM-based Constraint Mining, Constraints Test Generation, and Test Outcomes on RBCTest Dataset.
API Type Constraints Mining (RQ2) Test Gen. (RQ3) Test Out. (RQ4)
TP FP FN PRF1% N P×$\times$ ?
C.Holiday RP 16 0 8 100 66.7 80 16 16 100 12 0 4
G.Branch RR 36 2 13 94.7 73.5 82.8 36 35 97.2 33 2 0
G.Commit RR 55 2 18 96.5 75.3 84.6 55 54 98.2 40 8 6
G.Groups RR 63 2 22 96.9 74.1 84 62 61 98.4 60 1 0
G.Issues RR 93 2 48 97.9 66 78.8 93 92 98.9 80 7 5
G.Project RR 110 11 34 90.9 76.4 83 110 110 100 102 0 8
G.Repo. RR 33 3 11 91.7 75 82.5 33 33 100 30 2 1
Stripe RR 12 1 7 92.3 63.2 75 12 12 100 7 0 5
RP 17 0 4 100 81 89.5 17 16 94.1 14 1 1
Total 435 23 165 95.7 72.4 82.2 434 429 98.8 378 21 30
@@
7.2.1. Methodology
In RQ1, we assess the entire process from the input of API specifications to the test generation. In this RQ2, we focus on RBCTest’s first component, Static Constraint Mining via LLMs. Thus, we did not run AGORA. We used the RBCTest dataset with the available API specifications. To decide the constraints’ correctness in natural language, we use the following rules:
7.2.2. Results
As seen in Table 4, 95.7% of the constraints identified by our LLM-based static method are valid, although it missed 165 out of 600 constraints noted in the ground truth. Results for all metrics are notably higher than those of the entire process evaluated in RQ1, as in RQ1’s experiment, we aim to account for both the validity of constraints and the accuracy of the generated tests. As in Section 3, the detected constraints fall into two categories: Response Property constraints and Request-Response constraints. Our LLM-based static method’s constraint mining for Response Properties proves more effective than for Request-Response constraints, achieving 100% precision and 73.3% recall compared to 94.6% precision and 72.4% recall for Request-Response constraints.
Response Property constraints primarily define format and value ranges, with clear descriptions like “three-letter currency code” or “Unix epoch timestamp.” In contrast, Request-Response constraints are more complex, requiring precise mappings between request parameters and response properties. For example, in GitLab, ‘since’ and ‘until’ correctly map to ‘created_at’ to enforce range conditions. However, most false positives arise from incorrect mappings, especially in GitLab, where response body descriptions are lacking. Without details, mappings rely on attribute names, leading to errors—e.g., the ‘avatar’ parameter (for uploading images) is mistakenly linked to ‘avatar_url’ in the response due to name similarity.
@@ 7.3.1. Methodology
While RQ1 evaluates the entire process from constraint mining to test generation, this RQ3 focuses only on evaluating the test generation component for the constraints correctly detected by RBCTest via our LLM-based mining on the RBCTest dataset (thus, we did not run AGORA). We performed test generation from those correct constraints. The generated test cases are manually evaluated if they correctly verify the associated constraints.
We used the following rules to check if a test case is correct:
(1) Test Input: The generated test case is correct if it correctly receives two inputs for Request-Response constraints: 1) requested information and 2) response data, and one input for Response Property constraints: response data.
@@
(3) Test Output: The test must return:
@@
We only consider the set of test cases derived from valid constraints as identified in RQ2 (denoted as $N$ in Table 4).
@@ 7.3.2. Results
As shown in Table  4, RBCTest generated 434 test cases for 435 TP constraints (one test case was discarded by the verifier), including 401 for Request-Response constraints and 33 for Response Property constraints. It achieves a precision of 98.8% across 8 services and confirmed 429 correct test cases. Examining services, 4/8 services achieved a precision of 100%, while the lowest precision was in the Stripe service, with a precision of 94.1%.
Results indicate that RBCTest excels in generating code for Response Property constraints, which mainly involve format or value range validation. In contrast, Request-Response constraints are more error-prone due to the complex logic required to parse and verify dependencies between request parameters and response properties. Further analysis reveals that test generation errors primarily stem from (1) missing descriptions and (2) ambiguous keywords.
@@
7.4. Usefulness: Detecting Mismatches between Constraints and Response Bodies (RQ4)
@@ 7.4.1. Methodology
We use the generated test cases to detect mismatches between the constraints in the API specification and the API response bodies. With this goal, we did not run AGORA. For each API endpoint, we execute the API with multiple sets of request parameters. For each API execution, we collect 1) request information (i.e., what is expected from the API call), and 2) response data (i.e., the actual response). Response data contains multiple properties, each attached with constraints and generated constraint test cases. We ran the test cases associated with the response data to collect the outcomes. We also ran the inputs used in AGORA to verify against our test cases of correctly generated constraints. If the result is false, we report a mismatch. Otherwise, it is a match.
@@ 7.4.2. Results
We performed test runs on 429 correctly generated tests from RQ3 (Table 4). Our test results indicate 378 ‘matched’ response bodies (i.e., consistent with the specification), 21 ‘mismatched’, and 30 ‘unknown’. Out of 434 correctly mined constraints, 378 were verified by the tests, meaning that 87.1% of the constraints are met by the actual execution of the SUTs. Our tool detected 21 mismatches, revealing inconsistencies between specifications and execution of the APIs. An ‘unknown’ occurs when a property is absent in the response body due to its optional nature.
7.4.3. Analysis
@@
We revealed the following root causes of these mismatches: (1) incompatible data formats, (2) not-explicitly-described nullable properties, and (3) inter-parameter request dependencies.
@@
(1) Incompatible Data Formats: Our tool detected 21 constraint mismatches, mainly from GitLab services. For instance, the constraint ”date will be returned in ISO 8601 format YYYY-MM-DDTHH:MM" appears in GitLab operations. However, the actual data format is "2012-09-20T08:50:22.000Z", which includes a decimal part for seconds, leading to an inconsistency. Interestingly, we found that three instances of this type of inconsistency were reported as the issues on the GitLab forum (git, 2024c, d, a). This is anecdotal evidence on RBCTest’s usefulness in detecting real-world issues.
@@
(2) Not-Explicitly-Described, Nullable Properties: This issue is common in GitLab, where some response properties are nullable but lack descriptions in the specification, leading to parsing errors. Notably, these issues have been reported on the GitLab forum (git, 2024b).
@@
(3) Inter-Parameter Request Dependencies: This was found in only one case. For the operation ’GET/groups’ from GitLab Group, there is a constraint on the request parameter "order_by" affecting the "name" property in the response data. This logic dictates that the array of groups in the API response should be sorted according to the "order_by" parameter. RBCTest checked only one-to-one conditions among parameters, whereas the sorting order depends on both "order_by" and "sort" parameters (specifying the sort direction). As a result, this resulted in a detected mismatch.

### v3

@@ 8.2.1. Methodology
Using the best-performing result from Section 8.1, we generated validation scripts for all the mined constraints by following the process in Figure 7. We manually evaluated whether each verification script correctly reflects the intended logic described by its corresponding constraint. The evaluation followed the criteria mentioned during the construction of the ground truth as in Section 7. Verification scripts were executed against real-world API responses. Each constraint was verified by 50 API responses.
8.2.2. Results
For the AGORA/AGORA+ dataset, as seen in Table 6, RBCTest performs well in generating test scripts from mined constraints, achieving a validity rate of 86.4% and covering 76.6% of the ground-truth constraints. However, it missed 135 out of 578 constraints. Notably, for some services, e.g., Github_createOrgRepo and Github_getOrgRepo, RBCTest achieved up to 82% recall, while for Yelp, it reached 100% recall.
For the RBCTest dataset, Table 7 shows a precision of 91.7%, indicating that most generated test scripts correctly validate their associated constraints. The recall reached 71.3%, though 279 out of 973 constraints were missed. These metrics are slightly lower than those in RQ1, as this evaluation jointly considers both constraint correctness from the mining and the correctness of test execution.
RBCTest excels in test generation for Response Property constraints, which often involve clear format or range-based rules, e.g., “three-letter currency code” or “Unix epoch timestamp.” These constraints are easier to detect due to their descriptive nature. In contrast, Request–Response, Range-of-Value, and Computed Value constraints are more error-prone, as they require accurate mappings between request parameters and response properties. However, RBCTest still achieves very high precision on these constraints.
@@ 8.2.2. Results
Table 6. Constraint Test Generation (RQ2), Test Generation from Correct Constraints (RQ3), and Test Outcomes (RQ4) on the AGORA/AGORA+ Dataset.
API Constraint Test Gen (RQ2) Test Gen (RQ3) (Correct Constraints) Outcome (RQ4)
TP FP FN P% R% F1% N P% $\times$ ?
G.C 78 12 17 86.7 82.1 84.3 79 78 98.7 77 0 1
G.G 127 8 27 94.1 82.5 87.9 127 127 100.0 127 0 0
A.M 23 15 22 60.5 51.1 55.4 29 23 79.3 16 2 5
M.C 21 9 18 70.0 53.8 60.9 26 21 80.8 13 0 8
O.I 12 0 3 100.0 80.0 88.9 12 12 100.0 8 4 0
O.S 5 2 1 71.4 83.3 76.9 6 5 83.3 5 0 0
S.C 17 7 8 70.8 68.0 69.4 19 17 89.5 15 0 2
S.T 19 0 6 100.0 76.0 86.4 19 19 100.0 10 0 9
S.A 15 3 2 83.3 88.2 85.7 16 15 93.8 8 1 6
Y.B 4 1 0 80.0 100.0 88.9 4 4 100.0 3 1 0
Y.V 122 13 31 90.4 79.7 84.7 124 122 98.4 79 1 42
Total 443 70 135 86.4 76.6 81.2 461 443 96.1 361 9 73
@@
Table 7. Constraint Test Generation (RQ2), Test Generation from Correct Constraints (RQ3), Test Outcomes (RQ4) on RBCTest Dataset.
API Constraint Test Gen (RQ2) Test Gen (RQ3) (Correct Constraints) Outcome (RQ4)
TP FP FN P% R% F1% N P% $\times$ ?
C.H 36 0 6 100 85.7 92.3 36 36 100.0 22 2 12
G.B 47 6 13 88.2 75.0 84.7 47 45 95.7 42 1 2
G.C 64 7 16 90.1 75.3 88.5 69 64 92.8 51 6 7
G.G 89 6 26 93.7 76.1 85.8 91 89 97.8 78 6 4
G.I 171 8 91 95.5 64.3 78.7 175 171 97.7 132 19 20
G.P 163 16 70 91.1 68.5 80.6 168 163 97.0 139 0 24
G.R 44 7 11 86.3 80.0 83.0 46 44 95.7 41 0 2
S 82 13 26 86.3 74.5 82.0 84 82 97.6 68 3 11
Total 694 63 259 91.7 71.3 82.5 716 694 96.9 573 37 82
  • For test outcomes: (Matched), $\times$ (Mismatched), and ? (Unknown).
Test generation errors often stem from missing or vague descriptions, especially in services like GitLab, where response property documentation is sparse. For example, the request parameter avatar (used for image uploads) was incorrectly mapped to the response field avatar_url due to name similarity, despite different semantics. Services that define complex Data Format constraints—such as “numberOfNights must always be greater than 0,” or fields expected to be unique like id, code or barCode—also present challenges when descriptions are missing or implicit. Attributes with special rules (Name-based constraint) like latitude or longitude in AmadeusHotel_getMultiHotel or lang in Marvel were either overlooked or led to test scripts that failed to express the intended constraint logic. In these cases, the lack of structured or informative descriptions in the specification hindered constraint detection and test generation.
@@ 8.3.1. Methodology
The experiment in RQ2 evaluates the entire process from constraint mining to test generation. This RQ3 focuses only on evaluating the test generation for the constraints correctly mined by RBCTest. We perform test generation from those correct constraints. The generated test cases are manually evaluated if they correctly verify the associated constraints. We used the following rules to check if a test case is correct:
(1) Test Input: The generated test case is correct if it correctly receives two inputs for Request-Response constraints: 1) requested information and 2) response data, and one input for Response Property constraints (response data).
@@
(3) Test Output: Each script must be evaluated on 50 example responses collected from real-world data. The test must return:
@@
The final result is considered matched only if all test cases return 1 and none return -1; any case with 0 is acceptable. If all cases return 0, the result is unknown. If any case returns -1, the result is considered as mismatched. We only consider the set of test cases derived from valid constraints as identified in RQ2 (denoted as $N$ in Tables 67).
@@ 8.3.2. Results
As seen in Table 7, RBCTest generates 714 TP test cases out of 746 in RBCTest dataset, resulting a precision of 94.3%. In 4/8 services, it achieves a precision of 96%, while the lowest precision is in GitLab Repository API, with a precision of 83%. For the AGORA/AGORA+ dataset, it generates 443 correct test cases out of 513, with a precision of 96.1%. In 2 services, it achieves 100%.
Note that in RQ2, we consider all mined constraints, while in RQ3, we use only the correctly mined ones. Thus, the results in RQ3 (e.g., precision of 96.1% on the AGORA dataset) are higher than those in RQ2 (e.g., precision of 86.4%).
Our analysis reveals that RBCTest excels in generating code for Response Property constraints, which mainly involve format or value range validation. In contrast, Request-Response constraints are more error-prone due to the complex logic required to parse and verify dependencies between request parameters and response properties. Further analysis reveals that test generation errors primarily stem from (1) missing descriptions and (2) ambiguous keywords.
@@
8.4. Usefulness: Detecting Mismatches between Constraints and API Responses (RQ4)
@@ 8.4.1. Methodology
In this experiment, we use RBCTest’s generated test cases to detect mismatches between the API specification and API response bodies. A mismatch often indicates either that 1) the specification is outdated; or 2) the APIs’ implementation does not align with the constraints in their specification. To this goal, we performed the following. For the RBCTest dataset, for each API endpoint, we used RESTest (MartinLopez2021Restest) to generate 50 API requests, executed the SUT with APIs, and obtained the actual responses. For each execution, we collected 1) request data, and 2) actual response data. We used only the pairs of request-response with the status code of 2xx. Response data contains multiple properties, each attached with constraints and their test cases generated by RBCTest. We ran the test cases (in RQ3) generated for the constraints on the properties in the response data to collect the outcomes. If the result is false, we report a mismatch. Otherwise, it is a match.
For the AGORA/AGORA+ dataset, we used the request-response pairs provided on its repository, whose responses were obtained by its authors from the actual execution. We ran the test cases generated by RBCTest and compared with the responses as described.
@@ 8.4.2. Results
For RBCTest dataset, we performed test runs on all 694 correctly generated test cases from RQ3 (Table 7). The tests evaluate if the response data conforms to the constraints detected from the API specification. Our test results indicate 573 ‘matched’ response bodies (i.e., consistent with the specification), 37 ‘mismatched’, and 82 ‘unknown’. Out of 714 correctly mined constraints, 573 were verified by the test cases, meaning that 82.5% of the constraints are met by the actual execution of the SUTs. Our tool detected 37 mismatches, revealing inconsistencies between specifications and execution of the APIs. An ‘unknown’ occurs when a property is absent in the response body due to its optional nature.
For AGORA/AGORA+ dataset, we performed test runs on all 443 correctly generated tests from RQ3 (Table 6). Our test results indicate 361 ‘matched’ response bodies (i.e., consistent with the specification), 9 ‘mismatched’, and 73 ‘unknown’. Out of 443 correctly mined constraints, 361 were verified by the test cases, meaning that 81.5% of the constraints are satisfied by the actual execution of the APIs. Our tool detected 9 mismatches, revealing inconsistencies between specifications and API executions.
8.4.3. Analysis on Detected Mismatches
For 46 mismatches detected by RBCTest (detected-bugs), we conducted a detailed analysis by classifying them into 6 constraint categories, as illustrated in Figure  8. The isDateTime category, which refers to the constraint in date and time data types, accounts for the largest proportion (45.7%), indicating that temporal inconsistencies—such as variations in date–time formats or missing temporal fields—are the predominant mismatch between API specifications and their actual responses.
Figure 8. Detected Mismatches by Category
Other categories, including IsUrl (the property as a URL) and IsNumber (each 13.0%), IsBoolean and boolean values (10.9%), and Template Literals (e.g., String templates/lengths: Canadian provincial abbreviations in Canadian_Holidays API), and Value-In-Set (8.7%, e.g., in YouTube, video length in ISO_8601 duration formats: PT#M#S), show lower but still notable occurrences.
These findings suggest that the detected mismatches are non-trivial, and they mainly arise in the fields governed by strict formatting or parsing rules. In particular, the prevalence of date–time errors highlights the need for better normalization and schema-level validation. Beyond format-related issues, several mismatches may also stem from incorrect implementation or schema drift, where API implementations change faster than their specifications. This result indicates that RBCTest is useful in continuous validation and automated schema synchronization and in maintaining consistency between API behavior and its declared constraints.
8.4.4. Root-cause Analysis
@@
We reported the root causes of these mismatches: (1) incompatible data formats, (2) implicitly-described nullable properties, and (3) inter-parameter request dependencies.
@@
(1) Incompatible Data Formats: Our tool detected constraint mismatches mainly from GitLab services. For instance, the constraint ”date will be returned in ISO 8601 format YYYY-MM-DDTHH:MM" appears in GitLab operations. However, the actual data format is "2012-09-20T08:50:22.000Z", which includes a decimal part for seconds, leading to an inconsistency. Interestingly, we found that three instances of this type of inconsistency were reported as the issues on the GitLab forum (gitlab_issue_time1; gitlab_issue_time2; gitlab_issue_time3). This is anecdotal evidence on RBCTest’s usefulness in detecting real-world data-format issues.
@@
(2) Implicitly-Described, Nullable Properties: This issue is common in GitLab, where some response properties are nullable but lack descriptions in the specification, leading to parsing errors. Notably, these issues have been reported on the GitLab forum (gitlab).
@@
(3) Inter-Parameter Request Dependencies: This was found in only one case. For the operation ’GET/groups’ from GitLab Group, there is a constraint on the request parameter "order_by" affecting the "name" property in the response data. This logic dictates that the array of groups in the API response should be sorted according to the "order_by" parameter. RBCTest checked only one-to-one conditions among parameters, whereas the sorting order depends on both "order_by" and "sort" parameters (specifying the sort direction). As a result, this resulted in a detected mismatch. Future work could address more complex dependencies among multiple parameters.